OpenStudy (anonymous):

Can anyone help me out with this problem: A ball is thrown vertically upward with initial velocity v. Find the maximum height H of the ball as a function of v. Then find the initial velocity v required to achieve a height of H.

OpenStudy (anonymous):

we'll just use the eq, \(v^2 = u^2 +2aS\), since u=0, a=-g,(since it is opposite direction of v), rearranging gives \(S=\frac{v^2}{2g}\)

OpenStudy (anonymous):

where S=H

OpenStudy (anonymous):

ive arrived at this answer too...however when i submitted it...it said that g was not defined in this context

OpenStudy (anonymous):

is there any way to answer this in terms of H and v only?

OpenStudy (anonymous):

or, you can just sub g=10 into it

OpenStudy (anonymous):

thats what i would think so too. but it still says that answer is incorrect

OpenStudy (anonymous):

oh oh. sorry. my mistake. it's initial velocity, not end velocity. since v=0, \(u=\sqrt{2gH}=\sqrt{19.8H}\)

OpenStudy (anonymous):

tried that one too unfortunately....im not sure what they looking for

OpenStudy (anonymous):

dang==then i have no idea what they want

OpenStudy (anonymous):

thanks for trying to help though. ...its very much appreciated :)

OpenStudy (anonymous):

you're welcome :)